15/x - 15/x-2 = -2
My textbook tells me the answer is -3, 5 but I don't know how they got this.
Thank you for your help.

Answer by stanbon(48546)

Multiply both sides by LCD x(x-2) to get:
15(x-2)-15x = -2x(x-2)
-30 = -2x^2+4x
x^2-2x-15=0
(x-5)(x+3)=0
x=5 or x= -3
Cheers,
Stan H.